# Using the rejection sampling for finding tests

### v1

@@ 2.4 Testing if a population mean vector equals specific fixed vector
In this example, $p=2$. Two powerful tests are selected for comparison: the LR test and the empirical LR test (EL) (see, e.g., Owen,, 2001). Their rejection regions are determined using the asymptotic distribution $\chi^{2}_{2}$. The EL test is computed with the R package mvhtests. The statistical power of the AR test, LR test, and EL is estimated as the ratio of significant tests among 10000 Monte Carlo samples. The empirical power of these tests is illustrated in Figure 4.
@@ 2.5 A test for goodness-of-fit
Here, the goodness-of-fit type AR test described in Definition 2 is employed to test the hypothesis described above. Let $\widehat{f}$ denote a density estimate constructed, for example, using the vine copula-based kernel density estimator of Nagler and Czado, (2016), which is considered an efficient method for kernel density estimation in high-dimensional data. Here, the R package kdevine is used to compute the density estimates $\widehat{f}$. The test statistic $T(\textbf{X})$ is defined as the averaged sum of indicator variables depending on ratio between $f_{0}(X_{i})$ and $\widehat{f}(X_{i})$, evaluated at the sample values,
@@
The expected value of (6) can be compute applying the following Theorem (see also Kuismin,, 2025):
@@
$E\left[T(\textbf{X})\right]=n^{-1}\left[\#S_{1}+\sum_{s_{2}\in S_{2}}s_{2}\right].$ (7)
@@
$E\left[T(\textbf{X})\right]$ $=E\Bigl{\{}n^{-1}\sum_{i=1}^{n}\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]\Bigr{\}}=n^{-1}\sum_{i=1}^{n}P\Bigl{[}f_{0}(X_{i})/\widehat{f}(X_{i})>U\Bigr{]}$
@@
The AR test is compared with the Kolmogorov–Smirnov (KS) test (see, e.g., Giovanni and Franceschini,, 1987; Puritz et al.,, 2023), Cramér–von Mises (CVM) test (see, e.g., Chiu and Liu,, 2009), and Anderson-Darling test (AD) (Anderson and Darling,, 1952). While testing normality, a novel test proposed by Székely and Rizzo, (2005) (energy test (E)) is used. The AD and the energy tests are computed with R packages goftest and energy, respectively.
@@
The results summarized in Figure 5 show that the power of the AR test is higher than that of the KS and CVM tests, even for the small sample sizes considered in this example. For $n=20$ and $n=30$, the power of the AD test is slightly higher than that of the AR test when $\sigma<\sigma_{0}$. However, as the sample size increases, the AR test appears to be the most powerful among those considered in this example.
@@
When the AR test is applied as a test of multivariate normality, it is compared with five state-of-the-art multivariate procedures: the energy test, the test proposed by Zhou and Shao, (2014), referred to here as the $T_{n}$ test, test proposed by Henze and Zirkler (HZ test) (Henze and Zirkler,, 1990), Royston’s H test (Royston,, 1982), and test proposed by Villasenor-Alva and González-Estrada (VA-GE test) (Alva and Estrada,, 2009). The $T_{n}$ test, HZ test, Royston’s H test, and VA-GE test are computed with the R package mvnormalTest. The multivariate counterparts of the univariate distributions considered in the previous examples are used with dimension $p=3$. The empirical power of each test is defined as the ratio of significant tests over 1000 Monte Carlo samples at $\alpha=0.05$. Estimates of the power for the multivariate normality test are reported in Table 3.
@@
$t(2,\textbf{0},I)$ 0.05 0.84 0.82 0.78 0.83 0.64
$0.5N(\textbf{0},I)+0.5N(\textbf{3},I)$ 0.64 0.13 0.09 0.18 0.19 0.03
logistic(0, 1) 0.32 0.19 0.22 0.11 0.23 0.22
@@
$t(2,\textbf{0},I)$ 0.17 0.94 0.95 0.92 0.93 0.84
$0.5N(\textbf{0},I)+0.5N(\textbf{3},I)$ 0.93 0.25 0.16 0.40 0.35 0.04
logistic(0, 1) 0.81 0.24 0.30 0.16 0.26 0.29
@@
The results presented in Table 3 indicate that the power of the AR test is lower than that of the other tests when the alternative distribution is the multivariate $t$-distribution, although its power increases sharply (more than triples) as the sample size grows. In all other cases, the AR test is found to be the most powerful among the tests of multivariate normality considered.
@@ 3.2 Reaction times
Reaction time (RT), typically measured in milliseconds as the time required to complete a task, is a widely used dependent variable in psychology (see, e.g., Whelan,, 2008; Boeck and Jeon,, 2019). RT distributions are almost always right-skewed, and thus asymmetric, heavy-tailed distributions are commonly employed to capture their shape. In this example, the AR goodness-of-fit test (8) is applied to RT data from Experiment 1 of Wagenmakers et al., (2008). In this experiment, participants were instructed to respond either as quickly or as accurately as possible. Here, the analysis focuses on the distribution of RTs from the “quick” responses.
Because a detailed analysis of the RTs does not contribute directly to this work, only the RTs from the participant with ID 1 are analyzed. Censored samples (uninterpretable responses, RT <180<180 ms, or RT >3>3 sec) and duplicated values were removed prior to analysis ($n=366$). A visual inspection of the data suggests that a shifted log-normal distribution provides a good fit to the reaction times. Figure 6 displays the histogram of the data along with the fitted shifted log-normal distribution, where the mean, standard deviation, and shift parameters were estimated using the R package EnvStats.
Figure 6: Histogram of the RTs for participant ID 1, along with the fitted shifted log-normal distribution (solid red line).
The value of the AR test statistic is $0.961$ under $H_{0}$ that the RTs follow a shifted log-normal distribution with the estimated parameter values. The corresponding Monte Carlo $p$-value is approximately $0.894$. The $95\%$ credible interval for the statistic $T(\textbf{X})$, computed from the percentiles of the Poisson binomial distribution (see (3)) is $[0.940;0.978]$. Taken together, these results suggest that the shifted log-normal distribution provides a close fit to the RTs considered in this example.
@@ 4 Conclusion
Conflict of interest
The author declares no conflict of interest.

### v2

@@ 2.5 Testing if a population mean vector equals specific fixed vector
In this example, $p=2$. Two powerful tests are selected for comparison: the LR test and the empirical LR test (EL) (see, e.g., Owen, 2001). Their rejection regions are determined using the asymptotic distribution $\chi^{2}_{2}$. The EL test is computed with the R package mvhtests. The statistical power of the AR test, LR test, and EL is estimated as the ratio of significant tests among 10000 Monte Carlo samples. The empirical power of these tests is illustrated in Figure 4.
@@ 2.6 A test for goodness-of-fit
Here, the goodness-of-fit type AR test described in Definition 2 is employed to test the hypothesis described above. Let $\widehat{f}$ denote a density estimate constructed, for example, using kernel density estimation. Here, the R package Rfast2 using a high-degree memory and computationally optimized kernel functions is used to compute the density estimates $\widehat{f}$ Tsagris and Papadakis (2018). The test statistic $T(\textbf{X})$ is defined as the averaged sum of indicator variables depending on ratio between $f_{0}(X_{i})$ and $\widehat{f}(X_{i})$, evaluated at the sample values,
@@
The expected value of (6) can be compute applying the following Theorem (see also Kuismin, 2025):
@@
$E_{U}\left[T(\textbf{X})\right]=n^{-1}\left[\#S_{1}+\sum_{s_{2}\in S_{2}}s_{2}\right].$ (7)
@@
$E_{U}\left[T(\textbf{X})\right]$ $=E_{U}\Bigl\{n^{-1}\sum_{i=1}^{n}\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]\Bigr\}=n^{-1}\sum_{i=1}^{n}P\Bigl[f_{0}(X_{i})/\widehat{f}(X_{i})>U\Bigr]$
@@
The AR test is compared with the Kolmogorov–Smirnov (KS) test (see, e.g., Giovanni and Franceschini, 1987; Puritz et al., 2023), Cramér–von Mises (CVM) test (see, e.g., Chiu and Liu, 2009), and Anderson-Darling test (AD) (Anderson and Darling, 1952). While testing normality, a novel test proposed by Székely and Rizzo (2005) (energy test (E)) is used. The AD and the energy tests are computed with R packages goftest and energy, respectively.
@@
The results summarized in Figure 5 show that the power of the AR test is higher than that of the KS and CVM tests, even for the small sample sizes considered in this example. The power of the AR test is slightly higher than that of the AD test when $\sigma<\sigma_{0}$. As the sample size increases, the AR test appears to be the most powerful among those considered in this example.
@@
When the AR test is applied as a test of multivariate normality, it is compared with five state-of-the-art multivariate procedures: the energy test, the test proposed by Zhou and Shao (2014), referred to here as the $T_{n}$ test, test proposed by Henze and Zirkler (HZ test) (Henze and Zirkler, 1990), Royston’s H test (Royston, 1982), and test proposed by Villasenor-Alva and González-Estrada (VA-GE test) (Alva and Estrada, 2009). The $T_{n}$ test, HZ test, Royston’s H test, and VA-GE test are computed with the R package mvnormalTest. The multivariate counterparts of the univariate distributions considered in the previous examples are used with dimension $p=3$. The empirical power of each test is defined as the ratio of significant tests over 1000 Monte Carlo samples at $\alpha=0.05$. Estimates of the power for the multivariate normality test are reported in Table 3.
@@
$t(2,\textbf{0},I)$ 0.11 0.84 0.82 0.78 0.83 0.64
$0.5N(\textbf{0},I)+0.5N(\textbf{3},I)$ 0.99 0.13 0.09 0.18 0.19 0.03
logistic(0, 1) 0.76 0.19 0.22 0.11 0.23 0.22
@@
$t(2,\textbf{0},I)$ 0.09 0.94 0.95 0.92 0.93 0.84
$0.5N(\textbf{0},I)+0.5N(\textbf{3},I)$ 1.00 0.25 0.16 0.40 0.35 0.04
logistic(0, 1) 0.90 0.24 0.30 0.16 0.26 0.29
@@
The results presented in Table 3 indicate that the power of the AR test is lower than that of the other tests when the alternative distribution is the multivariate $t$-distribution. In all other cases, the AR test is found to be the most powerful among the tests of multivariate normality considered.
All of these examples show that the AR test maintains the Type I error at a pre-specified level of significance. The power of the AR test depends on the density estimation method. For example, instead of using the kernel density estimator from the R package Rfast2, one viable option is to employ the vine-copula-based kernel density estimator of Nagler and Czado (2016), which is considered an efficient method for kernel density estimation in high-dimensional settings. Using this estimator improves the power of the AR test when the alternative distribution is a heavy-tailed t-distribution. However, in the other cases considered here, the statistical power decreases slightly (results not shown).
@@ 3.2 Reaction times
Reaction time (RT), typically measured in milliseconds as the time required to complete a task, is a widely used dependent variable in psychology (see, e.g., Whelan, 2008; Boeck and Jeon, 2019). RT distributions are almost always right-skewed, and thus asymmetric, heavy-tailed distributions are commonly employed to capture their shape. In this example, the AR goodness-of-fit test (8) is applied to RT data from Experiment 1 of Wagenmakers et al. (2008). In this experiment, participants were instructed to respond either as quickly or as accurately as possible. Here, the analysis focuses on the distribution of RTs from the “quick” responses.
Because a detailed analysis of the RTs does not contribute directly to this work, only the RTs from the participant with ID 1 are analyzed. Censored samples (uninterpretable responses, RT <180<180 ms, or RT >3>3 sec) and duplicated values were removed prior to analysis ($n=366$). A visual inspection of the data suggests that a shifted log-normal distribution provides a good fit to the reaction times. Figure 6 displays the histogram of the data along with the fitted shifted log-normal distribution, where the mean, standard deviation, and shift parameters were estimated using the R package EnvStats, as well as the fitted normal distribution, for which the mean and standard deviation were estimated from the data.
Figure 6: Histogram of the RTs for participant ID 1, along with the fitted shifted log-normal distribution (solid red line), and normal distribution (dashed blue line).
The value of the AR test statistic is $0.961$ under $H_{0}$ that the RTs follow a shifted log-normal distribution with the estimated parameter values. The AR test statistic is $0.758$ under $H_{0}$ that the RTs follow a normal distribution. The corresponding Monte Carlo $p$-value is approximately $0.894$ and $0.001$, respectively. The $95\%$ credible interval for the statistic $T(\textbf{X})$, computed from the percentiles of the Poisson binomial distribution (see (3)) is $[0.940;0.978]$ and $[0.713;0.801]$, respectively. Taken together, these results suggest that the shifted log-normal distribution provides a close fit to the RTs considered in this example, offering a substantially better approximation than the normality assumption. Note that, although using estimated parameter values in hypothesis testing is questionable, this example illustrates how the AR test can be applied to evaluate different distributional assumptions rather than relying solely on the visual inspection of the data..